### v1

@@ 4 Experiments
In this section, we commence by presenting the experimental setups, and subsequently undertake a comprehensive evaluation of the nine prompting systems.
@@ 4.1 Experimental Setups
Datasets. We examine four visual attributes in the FR scenario, including synthetic structural and textural distortions, texture similarity, color differences, and geometric transformations. The selected datasets are FR-KADID [30], TQD [11], SPCD [51], and Aug-KADID [^2] We follow [11] to introduce mild geometric transformations. [30], respectively. In the NR scenario, we examine synthetic and authentic structural and textual distortions and algorithm-based artifacts, using the NR-KADID [30], SPAQ [16], and AGIQA-3K [25] datasets, respectively.
Sample Selection Details. In the FR scenario, the numbers of sampled reference images $N$ and the corresponding distorted images $K$ are set to $15$ and $10$, respectively. In the NR scenario, the number of selected images $N$ is set to $150$. The trade-off parameter $\lambda$ and the stability constant $\epsilon$ in Eqs. (3) and (5) are set to $0.01$ and $1$, respectively. The point-to-set diversity measure $\mathrm{Div}(\cdot)$ is implemented by the MSE in the feature space of the CLIP visual encoder [42]. The expert FR-IQA model $d_{w}(\cdot,\cdot)$ and the expert NR-IQA model $q_{w}(\cdot)$ are linearly fused with equal weightings using two top-performing models - one excelling in maintaining within-dataset correlation, and the other demonstrating robust cross-dataset generalization. Specifically, for within-dataset performance, we select TOPIQ [8] on FR-KADID [30], DISTS [11] on TQD [11] and Aug-KADID [30], CDNet [51] on SPCD [51], LIQE [71] on SPAQ [16] and NR-KADID [30], and Q-Align [57] on AGIQA-3K [25]. For cross-dataset performance, we choose LPIPS [69] and MANIQA [60] in the FR and NR scenarios, respectively.
Table 1: SRCC results of MLLMs paired with different prompting systems on uniformly sampled images. Model-S, Model-I, and Model-C denote the standard prompting, in-context prompting, and chain-of-thought prompting, respectively. The best result in each section is highlighted in bold.
@@ 4.2 Comparison of Nine Prompting Systems
We initially combine the nine prompting systems with three open-source MLLMs: LLaVA-v1.6 (Mistral-7B) [31], InternLM-XComposer2-VL (InernLM2-7B) [13] and mPLUG-Owl2 (LLaMA2-7B) [64] and one close-source MLLM: GPT-4V [61], and compare them on $150$ uniformly sampled images, with the goal of identifying the best prompting system for each MLLM. Table 1 shows the Spearman’s rank correlation coefficient (SRCC) results.
@@
Analysis of the Two IQA Scenarios. From the Table, it is evident that no open-source MLLMs achieve satisfactory IQA results in the FR scenario regardless of the prompting system used. These models may generate irrelevant text outputs or encounter complete failures when the input text prompts become more detailed and elaborate. We hypothesize that these models are predominantly trained or aligned on single-image vision tasks, making them struggle with analyzing multiple images, especially of the same underlying content [31]. Figure 5 provides a simple example of how different MLLMs behave when recognizing objects from multiple images, a high-level vision task at which they should excel. However, it appears that LLaVA-v1.6 [31] and mPLUG-Owl2 [64] completely disregard the second image, despite it being explicitly mentioned and separated from the first image in the input prompt. In the NR scenario with single-stimulus standard prompting (i.e., the single-image analysis scenario), LLaVA-v1.6 [31] and mPLUG-Owl2 [64] deliver improved quality prediction accuracy. Moreover, InternLM-XComposer2-VL [13] and GPT-4V perform remarkably in handling realistic camera distortions on SPAQ [16].
@@
Analysis of Psychophysical Prompting Methods. The results in the table reveal that for the three open-source MLLMs, the single-stimulus method is the optimal choice due to their limited ability to analyze multiple images. In stark contrast, GPT-4V [61] benefits from multiple-image analysis, and performs optimally under the double-stimulus chain-of-thought prompting. These findings raise questions about recent claims [58] that open-source MLLMs have surpassed GPT-4V-level IQA performance by instruction tuning on existing (weakly-)annotated benchmarks, given that the optimal prompting system has not been paired with GPT-4V.
@@
Figure 6: Comparison of the double-stimulus standard and chain-of-thought promptings for GPT-4V in the NR scenario.
@@
Analysis of NLP Prompting Methods. It is expected that in-context prompting is little likely to bring performance gains to the three open-source MLLMs due to the added complexity of processing additional image(s) in the context. Interestingly, we find that in-context prompting does not aid GPT-4V with the single-stimulus prompting. When combined with the multiple-stimulus prompting, GPT-4V performs poorly across various visual attributes and in both the FR and NR scenarios. To take a closer look, we compare the single-stimulus in-context prompting with double-stimulus standard prompting for GPT-4V. The main difference is that the former offers a human quality score for the contextual image, aiming to aid visual quality comparison. The comparable performance achieved by both prompting systems suggests that the inclusion of a human quality score has minimal impact on the GPT-4V inference. In cases when the input prompt contains a large number of images (e.g., eight images for the multiple-stimulus in-context prompting), even GPT-4V finds difficult in processing such substantial amount of visual information, resulting in a sharp drop in prediction accuracy.
@@
The chain-of-thought prompting, on the other hand, paints a different picture: it consistently enhances the performance of GPT-4V [61] in conjunction with three psychophysical testing protocols and across nearly all visual attributes. The remarkable improvements may arise because the chain-of-thought prompting elicits the break down of the intricate IQA task into simpler sub-tasks. This allows for a more meticulous inspection of visually critical factors such as structure and texture preservation, as well as color and luminance reproduction. On the other hand, as shown in Figure 6, the chain-of-thought prompting encourages absorbing IQA into a boarder and higher-level task, orchestrating with additional physical, geometrical, and semantic information of the natural scene to infer image quality. This reminds the authors of a famous slogan in computer vision: “If you cannot solve a simple problem in vision, you may have to solve a complex one (by Songchun Zhu)”.
@@
Table 2: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the FR scenario. Black and blue numbers in bold represent the best expert IQA systems and the MLLMs, respectively. * indicates that the model has been trained on the dataset.
@@ 4.3 Further Testing on Difficult Data
In this subsection, we compare the previous four MLLMs, each with the optimally suited prompting system and a quality-instruction-tuned MLLM, Q-Instruct [56], against representative expert IQA systems, including PSNR, SSIM [52], FSIM [68], LPIPS [69], AHIQ [24] and DISTS [11] as FR models, and NIQE [37], MUSIQ [21], MANIQA [60], LIQE [71] as NR models on the set of difficult data described in Sec. 3.3.
Results in the FR Scenario. Table 2 shows the SRCC and Pearson’s linear correlation coefficient (PLCC) results in the FR scenario. The primary observation is that the selected difficult images by “attacking” expert IQA models using Eqs. (3) and (5) pose a challenge to GPT-4V, with reduced IQA performance. Nevertheless, on FR-KADID with synthetic structural and textural distortions, GPT-4V [61] is on par with the leading expert IQA model, FSIM [68], which assesses structural similarity in the feature space. On Aug-KADID and TQD with geometric transformations and texture similarity, respectively, GPT-4V exhibits a commendable level of resilience, although not as good as DISTS [11]. It is important to recognize that the ways of attaining these aspects of perceptual robustness differ between the two methods. DISTS relies on comparison of spatial averages of VGG [45] feature maps as a complete and holistic texture characterization, whereas GPT-4V benefits from its intrinsic ability to compile and process data sequentially (as highlighted in the Introduction). On SPCD with color differences, all MLLMs including GPT-4V encounter challenges in emulating human color perception, and struggle to differentiate images even with clearly noticeable variations in color appearances (to the human eye).
@@
Results in the NR Scenario. Table 3 shows the SRCC and PLCC results in the NR scenario. It is clear that GPT-4V [61] is outstanding in capturing the authentic structural and textural distortions on SPAQ, surpassing the two expert IQA systems MUSIQ [21] and LIQE [71]. Notably, the latter has prior exposure to the SPAQ dataset. Furthermore, GPT-4V showcases a remarkable generalizablity to AI-generated images, indicating its great potential to guide the optimization of generative AI models for images. Q-Instruct [56], a model fine-tuned from a variant of LLaVA, exhibits significantly enhanced IQA skills compared to LLaVA-v1.6, thus highlighting the effectiveness of visual instruction tuning (particularly on images with similar distortion appearances).
@@ 4.4 Discussion and Limitation
We briefly discuss some of the limitations of this work and opportunities for future work. First, the current prompting systems have room for improvement, as the input prompts are not optimized. This presents an opportunity for exploring automatic prompt optimization [27, 18] within our prompting systems. Second, our sampler relies on human quality scores, and is thus only applicable to existing labeled IQA datasets. Extending it to sample from large-scale sets of unlabeled images will need to 1) accelerate the inference speed of MLLMs so as to leverage model falsification methodologies (e.g., the gMAD competition [36]) or 2) train additional failure-prediction modules [7] in a parameter-efficient way [19]. Third, the textual responses produced by MLLMs have not been quantitatively assessed. Prior research has investigated the application of GPT-4 [31] to assess these responses in terms of correctness, consistency, relevance, informativeness, coherence, and naturalness. Given our focus on responses pertinent to visual quality, it seems reasonable to develop a semantic textual similarity model, specifically designed for the IQA task, with improved accuracy and robustness. Fourth, this work does not touch on instruction tuning of MLLMs to enhance the IQA performance. Preliminary efforts have been made by Wu et al. [56] and You et al. [67] to directly fine-tune open-source MLLMs on datasets with image quality descriptions. Our results suggest a more encouraging approach: incorporating IQA in a boarder vision task for joint optimization, with the goal of achieving an ideal balance between the generality and specificity of pre-trained MLLMs [61, 33, 23].
@@ 5 Conclusion
We have presented a comprehensive study of MLLMs for IQA, with emphasis on systematic prompting strategies. Our study arises as a natural combination of methods from two separate lines of research: psychophysics and NLP. The first involves collecting reliable measurements of perceptual quantities in response to physical stimuli. The second endeavours to design input textual descriptions to an LLM to elicit a specific response. Our experiments have shown that different MLLMs admit different prompting systems to work optimally. This suggests the need for a systematic re-evaluation of the recent progress in MLLMs for IQA, particular in comparison to the state-of-the-art GPT-4V [61]. Moreover, there is still ample room in improving MLLMs (including GPT-4V) for IQA, especially fine-grained quality discrimination and multiple-image quality analysis.

### v2

@@ 4 Experiments
In this section, we first present the experimental setups, and then carry out a comprehensive evaluation of the nine IQA prompting systems for MLLMs.
@@ 4.1 Experimental Setups
Datasets. We examine four visual attributes in the FR scenario, including synthetic structural and textural distortions, geometric transformations, texture similarity, and color differences. The selected datasets are FR-KADID [28], TQD [11], SPCD [49], and Aug-KADID [^2] We follow [11] to introduce mild geometric transformations. [28], respectively. In the NR scenario, we examine synthetic and authentic structural and textual distortions and algorithm-based artifacts, using the NR-KADID [28], SPAQ [15], and AGIQA-3K [23] datasets, respectively.
Sample Selection Details. In the FR scenario, the numbers of sampled reference images $N$ and the corresponding distorted images $K$ are set to $15$ and $10$, respectively. In the NR scenario, the number of selected images $N$ is set to $150$. The trade-off parameter $\lambda$ and the stability constant $\epsilon$ in Eqs. (1) and (3) are set to $0.01$ and $1$, respectively. The point-to-set distance measure $\mathrm{Div}(\cdot)$ is implemented by the MSE in the feature space of the CLIP visual encoder [39]. The expert FR-IQA model $d_{w}(\cdot,\cdot)$ and NR-IQA model $q_{w}(\cdot)$ are linearly fused with equal weightings using two top-performing models - one excelling in maintaining within-dataset correlation, and the other demonstrating robust cross-dataset generalization. Specifically, for within-dataset performance, we select TOPIQ [8] on FR-KADID [28], DISTS [11] on TQD [11] and Aug-KADID [28], CDNet [49] on SPCD [49], LIQE [70] on SPAQ [15] and NR-KADID [28], and Q-Align [55] on AGIQA-3K [23]. For cross-dataset performance, we choose LPIPS [68] and MANIQA [58] in the FR and NR scenarios, respectively.
Table 1: SRCC results of MLLMs paired with different prompting systems on uniformly sampled images. Model-S, Model-C, and Model-I denote the standard prompting, chain-of-thought prompting, and in-context prompting, respectively. The best result in each section is highlighted in bold.
@@ 4.2 Comparison of Nine Prompting Systems
We initially combine the nine prompting systems with three open-source MLLMs: LLaVA-v1.6 (Mistral-7B) [29], InternLM-XComposer2-VL (InernLM2-7B) [13] and mPLUG-Owl2 (LLaMA2-7B) [62] and one closed-source MLLM: GPT-4V [59], and compare them on $150$ uniformly sampled images, with the goal of identifying the best prompting system for each MLLM. Table 1 shows the Spearman’s rank correlation coefficient (SRCC) results.
@@
Analysis of Two IQA Scenarios. From the Table, it is evident that no open-source MLLMs achieve satisfactory IQA performance in the FR scenario regardless of the adopted prompting system. These models may generate irrelevant text outputs or encounter complete failures when the input text prompts become more detailed and elaborate. We hypothesize that these models are predominantly trained or aligned on single-image vision tasks, making it challenging for them to analyze multiple images, especially of the same underlying content [29]. Figure 5 provides an example of how different MLLMs behave when recognizing objects from multiple images, a high-level vision task at which they should excel. However, it appears that LLaVA-v1.6 [29] and mPLUG-Owl2 [62] completely disregard the second image, despite it being explicitly mentioned and separated from the first image in the input prompt. In the NR scenario with single-stimulus standard prompting (i.e., the single-image analysis scenario), LLaVA-v1.6 [29] and mPLUG-Owl2 [62] deliver improved quality prediction accuracy. InternLM-XComposer2-VL [13] and GPT-4V perform remarkably in handling realistic camera distortions on SPAQ [15].
@@
Analysis of Psychophysical Prompting Methods. The results in the table reveal that for the three open-source MLLMs, the single-stimulus method is the optimal choice due to their limited ability to analyze multiple images. In stark contrast, GPT-4V [59] benefits from multiple-image analysis, and performs optimally under double-stimulus chain-of-thought prompting. These findings raise questions about recent claims [56, 66, 65] that open-source MLLMs have surpassed GPT-4V-level IQA performance by instruction-tuning on existing (weakly-)annotated benchmarks, given that the optimal prompting system has not been paired with GPT-4V.
@@
Figure 6: Comparison of double-stimulus standard and chain-of-thought promptings for GPT-4V in the NR scenario.
@@
Analysis of NLP Prompting Methods. It is expected that in-context prompting is little likely to bring performance gains to the three open-source MLLMs due to the added complexity of processing additional image(s) in the context. Interestingly, we find that in-context prompting does not aid GPT-4V with single-stimulus prompting either. When combined with multiple-stimulus prompting, GPT-4V performs poorly across various visual attributes and in both the FR and NR scenarios. Taking a closer look, we compare single-stimulus in-context prompting with double-stimulus standard prompting for GPT-4V. The main difference is that the former offers a human quality score for the contextual image, aiming to aid visual quality comparison. The comparable performance achieved by both prompting systems suggests that the inclusion of a human quality score has minimal impact on the GPT-4V inference. In cases when the input prompt contains multiple images (e.g., eight images for multiple-stimulus in-context prompting), even GPT-4V finds difficulty in processing such substantial amount of visual information, resulting in a sharp performance drop.
@@
The chain-of-thought prompting, on the other hand, paints a different picture: it consistently enhances the performance of GPT-4V [59] under three psychophysical testing protocols and across nearly all visual attributes. The remarkable improvements may arise because the chain-of-thought prompting elicits the break down of the intricate IQA task into simpler sub-tasks. This allows for a more meticulous inspection of visually critical factors such as structure and texture preservation, as well as color and luminance reproduction. On the other hand, as shown in Figure 6, the chain-of-thought prompting also encourages integrating IQA into a broader and higher-level task, orchestrating with additional physical, geometrical, and semantic information of the natural scene to infer image quality. This reminds the authors of a famous slogan in computer vision: “If you cannot solve a simple problem in vision, you may have to solve a complex one (by Songchun Zhu)”.
@@
Table 2: Comparison of MLLMs with optimally suited prompting systems against expert IQA systems in the FR scenario. * indicates that the model has been trained on the dataset.
@@ 4.3 Further Testing on Difficult Data
In this subsection, we compare the previous four MLLMs, each with the optimally suited prompting system and a quality-instruction-tuned MLLM, Q-Instruct [54], against representative expert IQA systems, including PSNR, SSIM [50], FSIM [67], LPIPS [68], AHIQ [22] and DISTS [11] as FR models, and NIQE [34], MUSIQ [20], MANIQA [58] and LIQE [70] as NR models on the set of difficult data described in Sec. 3.3.
Results in the FR Scenario. Table 2 shows the SRCC and Pearson’s linear correlation coefficient (PLCC) results in the FR scenario. The primary observation is that the selected difficult images by “attacking” expert IQA models using Eqs. (1) and (3) pose a challenge to GPT-4V, with noticeably reduced IQA performance. Nevertheless, on FR-KADID with synthetic structural and textural distortions, GPT-4V [59] is on par with the leading expert IQA model, FSIM [67], which assesses structural similarity in the feature space. On Aug-KADID and TQD with geometric transformations and texture similarity, respectively, GPT-4V exhibits a commendable level of resilience, although not as good as DISTS [11]. It is important to note that the ways of attaining such perceptual robustness differ between the two methods. DISTS relies on comparison of spatial averages of VGG [42] feature maps, whereas GPT-4V benefits from its intrinsic ability to compile and process data sequentially. On SPCD with color differences, all MLLMs, including GPT-4V, encounter challenges in emulating human color perception, and struggle to differentiate images even with clearly noticeable variations in color appearances.
@@
Results in the NR Scenario. Table 3 shows the SRCC and PLCC results in the NR scenario. It is clear that GPT-4V [59] is outstanding in capturing the authentic structural and textural distortions on SPAQ, surpassing the two expert IQA systems MUSIQ [20] and LIQE [70]. Furthermore, GPT-4V showcases a remarkable generalizablity to AI-generated images, indicating its great potential to guide the optimization of AI generative models for images. Q-Instruct [54], fine-tuned from a variant of LLaVA, exhibits enhanced IQA skills compared to LLaVA-v1.6, thus highlighting the effectiveness of visual instruction tuning.
@@ 4.4 Discussion and Limitation
We briefly discuss some of the limitations of this work and opportunities for future work. First, the current prompting systems have room for improvement, as the input prompts are not optimized. This presents an opportunity for exploring automatic prompt optimization [25, 17] within our prompting systems. Second, our sampler relies on human quality scores, and is thus only applicable to existing human-rated IQA datasets. Extending it to sample from large-scale unlabeled image sets will need to 1) accelerate the inference speed of MLLMs so as to leverage model falsification methodologies (e.g., the gMAD competition [33]) or 2) train additional failure-prediction modules [7] in a parameter-efficient way [18]. Third, the textual responses produced by MLLMs have not been quantitatively assessed. Prior research has applied GPT-4 [2] to assess quality responses in terms of correctness, consistency, relevance, informativeness, coherence, and naturalness. Fourth, this work does not touch on instruction tuning of MLLMs to enhance the IQA performance. Preliminary efforts have been made by Wu et al. [54] and You et al. [66] to directly fine-tune open-source MLLMs on datasets with image quality descriptions. Our results suggest a more encouraging approach: (active) continual learning and/or parameter-efficient tuning of MLLMs to strike a good balance between the specificity (to IQA) and the generality of open-source MLLMs.
@@ 5 Conclusion
We have presented a comprehensive study of MLLMs for IQA, with emphasis on systematic prompting strategies. Our study arises as a natural combination of methods from two separate lines of research: psychophysics and NLP. The first involves collecting reliable measurements of perceptual quantities in response to physical stimuli. The second endeavors to design input textual descriptions to an LLM to elicit a specific response. Our experiments have shown that different MLLMs admit different prompting systems to perform optimally. This suggests the need for a re-evaluation of the recent progress in MLLMs for IQA. Moreover, there is still ample room for improving the IQA capabilities of MLLMs (including GPT-4V), especially in terms of fine-grained quality discrimination and multiple-image quality analysis.